OpenStudy (anonymous):

how to find the sum from 0 to infinity (2^(K+3))/3^K

OpenStudy (anonymous):

\[\sum_{0}^{\infty} 2^{k+3}/3^{k}\]

OpenStudy (amistre64):

2^k 2^3/3^k

OpenStudy (amistre64):

got it

OpenStudy (amistre64):

\[\sum2^3(\frac{2}{3})^k\to 8\sum (\frac{2}{3})^k \]

OpenStudy (amistre64):

the sum part goes to 1/(1-2/3)

OpenStudy (amistre64):

1/(1/3) = 3 8*3 = 24

OpenStudy (anonymous):

great explanation! Got it man!

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

what if the denominator goes from 3^k to 3^(K+3) and then numerator became 1

OpenStudy (amistre64):

same concepts, just different numbers really

OpenStudy (anonymous):

\[\sum_{0}^{\infty} 1/3^{k+3}\]

OpenStudy (amistre64):

\[B^{a+b}\to B^aB^b\] that allow you to factor off the constant and use the rest as a ratio

OpenStudy (amistre64):

1/3^3 sum (1/3)^k

OpenStudy (amistre64):

the sum of a geometric sequence is:\[\sum \frac{1-r^n}{1-r}\to \frac{1-r^{inf}}{1-r}\] when r is less than 1, r^inf goes to zero

OpenStudy (amistre64):

well, when |r|<1 ...

OpenStudy (anonymous):

true now I'm struggling on this question \[\sum_{0}^{\infty} 3^{k-1}/4^{3k+1}\] i try to use your method, but didn't work I have no idea how to make the denominator good

OpenStudy (amistre64):

well, lets start by stripping off the boring things :)

OpenStudy (amistre64):

\[ \sum_{0}^{\infty} 3^{k-1}/4^{3k+1}\] \[ \sum_{0}^{\infty} 3^{k}3^{-1}/4^{3k}4^{1}\] \[ 1/12\sum_{0}^{\infty} 3^{k}/4^{3k} \] \[ 1/12\sum_{0}^{\infty} 3^{k}/(4^3)^k \] \[ 1/12\sum_{0}^{\infty} (3/64)^k \]

OpenStudy (amistre64):

now r=3/64

OpenStudy (anonymous):

oh ya!!